# History, Development, and Principles of Large Language Models—An Introductory Survey

### v1

@@ 4. APPLICATIONS OF LARGE LANGUAGE MODELS
Drug Discovery: The field of drug discovery is characterized by substantial challenges and elevated costs. The utilization of computational methodologies, such as AI, DL, and quantum mechanical techniques, holds the promise of accelerating the identification of potential drug candidates and predicting their properties (Sadybekov & Katritch, 2023; Gorgulla et al., 2022). Among recent advancements in this arena, the integration of LLMs, such as ChatGPT, emerges as a valuable addition, adept at generating and analyzing natural language texts (Savage, 2022). LLMs contribute significantly to the initial phase of drug discovery, particularly in identifying suitable drug targets for diverse diseases (Haley & Roudnicky, 2020; Paananen & Fortino, 2020). They offer preliminary insights into the structure and functionality of protein-based drug targets, subject to further validation by alternative methodologies. Furthermore, LLMs play a pivotal role in subsequent drug discovery steps, aiding in the design and screening of small molecules capable of interacting with these drug targets (Savage, 2022).
@@
Finance: LLMs are driving significant advancements in the finance industry, with significant financial applications including trading and portfolio management (Zhang et al., 2020), financial risk modeling (Mashrur et al., 2020), financial text mining (Gupta et al., 2020; Pagliaro et al., 2021), and financial advisory and customer services (Shah et al., 2020). AI integration in financial advisory and customer services is an emerging and rapidly expanding field. According to (Misischia et al., 2022), AI-driven chatbots already provide more than 37% of supporting functions in various e-service scenarios. Within the financial field, these chatbots are embraced as cost-effective alternatives to human customer service, a trend underscored in the ”Chatbots in Consumer Finance” report [^7] https://www.consumerfinance.gov/data-research/research-reports/chatbots-in-consumer-finance/chatbots-in-consumer-finance. With the advent of LLMs, more and more tasks that were previously considered difficult to handle have become possible, further expanding the potential applications of AI in the financial industry (Li et al., 2023). Notably, there are already some specialized LLMs like BloombergGPT (Wu et al., 2023), a 50-billion-parameter model excelling in financial tasks while maintaining strong performance in general language model benchmarks.
@@
Medical: LLMs are now at the forefront of medical AI with immense potential to improve the efficiency and effectiveness of clinical, educational, and research work (Thirunavukarasu et al., 2023). One example is the performance of LLMs on the MedQA dataset (Jin et al., 2021), which serves as a widely used biomedical question-answering dataset consisting of the United States Medical Licensing Examination (USMLE)-style questions. In this dataset, ChatGPT attains human passing levels and (kung et al., 2023). Additionally, in a span of fewer than six months, Med-PaLM2 achieves proficiency levels close to human experts (Singhal et al., 2023). The strong performance of GPT-4 (OpenAI, 2023) and Med-PaLM2 in medical tests suggests that LLMs may be useful teaching tools for students currently attaining a lower level in such tests (Nori et al., 2023). GPT-4’s meta-prompt feature allows users to explicitly describe the desired role for the chatbot to take on during conversation; a useful example is a “Socratic tutor mode” [^8] https://sites.highlands.edu/tutorial-center/tutor-resources/online-tutor-training/module-4/the-socratic-method/, which encourages students to think for themselves by pitching questions at decreasing levels of difficulty until students can work out solutions to the fuller question at hand. This mode can foster critical thinking and problem-solving skills among students (Thirunavukarasu et al., 2023). ChatGPT demonstrates superior performance in tasks not necessitating specialized knowledge or when such expertise is available in user prompts (Arora et al., 2023). For example, consider discharge summaries as a case in point - they entail repetitive tasks involving information interpretation and compression, often requiring minimal specialized knowledge and user prompts (Touvron et al., 2023).
@@
Legal: LLMs exhibit robust capabilities across various legal tasks, particularly in case prediction and generating legal texts. For instance, ChatGPT achieved a 50.3% accuracy rate on NCBE MBE practice exams, significantly surpassing the 25% random guess level, and achieved passing scores in both Evidence and Torts (Bommarito & Katz, 2022). The integration of ChatGPT holds the potential to reshape the entire legal industry. An Australian law firm experimented with ChatGPT to draft a statement of claim based on the 1992 Mabo case, yielding results akin to those of a first-year lawyer [^9] How ChatGPT and other new AI tools are being used by lawyers, architects and coders - ABC News . Iu and Wong (2022) investigation explored the possibility that ChatGPT could replace litigators by evaluating its drafting and research capabilities. While recommending ChatGPT as a complement to litigators rather than a complete replacement, this suggestion stemmed from ChatGPT’s limitations in meeting the precise accuracy and timeliness demands of legal texts, especially in ensuring continuous access to updated legal research. More recently, advancements such as DeLilegal [^10] https://data.delilegal.com/lawQuestion addressed this issue by integrating with external law databases to enhance retrieval and minimize hallucination. Meanwhile, LLMs in the legal field have emerged in recent months such as ChatLaw (Cui et al., 2022).
@@ 5. DRAWBACKS AND FUTURE DIRECTIONS OF LARGE LANGUAGE MODELS
Fairness: LLMs are widely employed for decision-making across various domains, impacting individuals and society. However, these decisions might exhibit bias against marginalized populations. For instance, when utilized in screening resumes for programming positions, LLMs might exhibit a bias favoring male candidates, demonstrating evident gender discrimination. This bias stems from the training of LLMs on extensive and unstructured Internet data, where they inherit stereotypes, misrepresentations, derogatory language, and exclusionary behaviors. These aspects disproportionately affect already disadvantaged and marginalized groups (Bender et al., 2021; Dodge et al., 2021; Sheng et al., 2021). The commonly employed strategy to mitigate bias in LLMs is to remove biased data from the training dataset. However, this strategy does not entirely eliminate bias and often diminishes the effectiveness of language modeling (Meade et al., 2021). Saxena et al. (2020) discuss that the root causes of these problems lie not only in technology but also in socially acceptable definitions of fairness and meaningful interventions to ensure the long-term well-being of all groups. Researchers should consider the needs of disadvantaged groups from the outset and design these technologies proactively, rather than simply reacting to the biases present in their designed systems (Gallegos et al., 2023). The documented bias in LLMs, as highlighted by studies like (Blodgett & O’Connor, 2017Mei et al., 2023Mozafari et al., 2020), underscores the imperative for future efforts in this direction.
@@
Safety: LLMs have diverse applications in industries such as finance and healthcare. However, these applications also present significant security challenges. LLMs may generate outputs that diverge from the provided context, user input, or factual knowledge, a phenomenon referred to as “hallucination” (Bangs et al., 2023; Huang et al., 2023). This undermines their reliability in real-world scenarios, emphasizing the vital necessity for precision, particularly in critical domains like medicine (Dash et al., 2023; Umapathi et al., 2023). As a primary way to avoid these problems, well-aligned LLMs can be developed by including humans in the training loop and using reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022; Glaese et al., 2023). To improve model safety, it is also important to include safety-related cues in the RLHF process, as shown in GPT-4 (OpenAI, 2023). However, RLHF relies heavily on high-quality human feedback data from professional annotators, which is costly and time-consuming. As a result, enhancing the RLHF framework is essential to alleviate the burden on human annotators. Additionally, exploring more efficient annotation methods with assured data quality, such as Large Language Models (LLMs), can assist in the annotation process and enhance overall safety.
@@
Intellectual Property: LLMs possess the capability to create content resembling human-generated material, potentially encroaching upon users’ intellectual property rights. These models rely on training and optimizing using users’ textual data, which may encompass personal information, proprietary knowledge, patented technology, and more. Notably, AI-generated code exemplifies this issue. Code-generation models trained on open-source repositories can produce programs resembling existing ones, potentially disregarding relevant licenses (Yu et al, 2023). For instance, legal action has been taken against Microsoft, GitHub, and OpenAI, alleging copyright infringement due to Copilot reproducing licensed code without adherence to licensing terms [^11] https://www.reuters.com/technology/microsoft-attracting-users-its-code-writing-generative-ai-software-2023-01-25/  [^12] https://www.reuters.com/legal/litigation/openai-microsoft-want-court-toss-lawsuit-accusing-them-abusing-open-source-code-2023-01-27/. Consequently, there’s a pressing need for heightened attention and regulation concerning intellectual property rights concerning LLMs. This is essential to safeguard the legitimate rights and interests of original creators, users, and the public while fostering and respecting the innovation and advancement of LLMs technology (Li et al., 2023).

### v2

@@ 4. APPLICATIONS OF LARGE LANGUAGE MODELS
Drug Discovery: The field of drug discovery is characterized by substantial challenges and elevated costs. The utilization of computational methodologies, such as AI, DL, and quantum mechanical techniques, holds the promise of accelerating the identification of potential drug candidates and predicting their properties [sadybekov2023computational, gorgulla2022emerging]. Among recent advancements in this arena, the integration of LLMs, such as ChatGPT, emerges as a valuable addition, adept at generating and analyzing natural language texts [savage2023drug]. LLMs contribute significantly to the initial phase of drug discovery, particularly in identifying suitable drug targets for diverse diseases [haley2020functional, paananen2020omics]. They offer preliminary insights into the structure and functionality of protein-based drug targets, subject to further validation by alternative methodologies. Furthermore, LLMs play a pivotal role in subsequent drug discovery steps, aiding in the design and screening of small molecules capable of interacting with these drug targets [savage2023drug].
@@
Finance: LLMs are driving significant advancements in the finance industry, with significant financial applications including trading and portfolio management [zhang2020deep], financial risk modeling [mashrur2020machine], financial text mining [gupta2020comprehensive, pagliaro2021investor], and financial advisory and customer services [shahfinaid]. AI integration in financial advisory and customer services is an emerging and rapidly expanding field. According to [misischia2022chatbots], AI-driven chatbots already provide more than 37% of supporting functions in various e-service scenarios. Within the financial field, these chatbots are embraced as cost-effective alternatives to human customer service, a trend underscored in the ”Chatbots in Consumer Finance” report [^7] https://www.consumerfinance.gov/data-research/research-reports/chatbots-in-consumer-finance/chatbots-in-consumer-finance. With the advent of LLMs, more and more tasks that were previously considered difficult to handle have become possible, further expanding the potential applications of AI in the financial industry [li2023large]. Notably, there are already some specialized LLMs like BloombergGPT [wu2023bloomberggpt], a 50-billion-parameter model excelling in financial tasks while maintaining strong performance in general language model benchmarks.
@@
Medical: LLMs are now at the forefront of medical AI with immense potential to improve the efficiency and effectiveness of clinical, educational, and research work [thirunavukarasu2023large]. One example is the performance of LLMs on the MedQA dataset [jin2021disease], which serves as a widely used biomedical question-answering dataset consisting of the United States Medical Licensing Examination (USMLE)-style questions. In this dataset, ChatGPT attains human passing levels and [kung2023performance]. Additionally, in a span of fewer than six months, Med-PaLM2 achieves proficiency levels close to human experts [singhal2023towards]. The strong performance of GPT-4 [achiam2023gpt] and Med-PaLM2 in medical tests suggests that LLMs may be useful teaching tools for students currently attaining a lower level in such tests [thirunavukarasu2023large]. GPT-4’s meta-prompt feature allows users to explicitly describe the desired role for the chatbot to take on during conversation; a useful example is a “Socratic tutor mode” [^8] https://sites.highlands.edu/tutorial-center/tutor-resources/online-tutor-training/module-4/the-socratic-method/, which encourages students to think for themselves by pitching questions at decreasing levels of difficulty until students can work out solutions to the fuller question at hand. This mode can foster critical thinking and problem-solving skills among students [thirunavukarasu2023large]. ChatGPT demonstrates superior performance in tasks not necessitating specialized knowledge or when such expertise is available in user prompts [arora2023promise]. For example, consider discharge summaries as a case in point - they entail repetitive tasks involving information interpretation and compression, often requiring minimal specialized knowledge and user prompts [touvron2023llama].
@@
Legal: LLMs exhibit robust capabilities across various legal tasks, particularly in case prediction and generating legal texts. For instance, ChatGPT achieved a 50.3% accuracy rate on NCBE MBE practice exams, significantly surpassing the 25% random guess level, and achieved passing scores in both Evidence and Torts [bommarito2022gpt]. The integration of ChatGPT holds the potential to reshape the entire legal industry. An Australian law firm experimented with ChatGPT to draft a statement of claim based on the 1992 Mabo case, yielding results akin to those of a first-year lawyer [^9] How ChatGPT and other new AI tools are being used by lawyers, architects and coders - ABC News . [iu2023chatgpt] investigation explored the possibility that ChatGPT could replace litigators by evaluating its drafting and research capabilities. While recommending ChatGPT as a complement to litigators rather than a complete replacement, this suggestion stemmed from ChatGPT’s limitations in meeting the precise accuracy and timeliness demands of legal texts, especially in ensuring continuous access to updated legal research. More recently, advancements such as DeLilegal [^10] https://data.delilegal.com/lawQuestion addressed this issue by integrating with external law databases to enhance retrieval and minimize hallucination. Meanwhile, LLMs in the legal field have emerged in recent months such as ChatLaw [cui2023chatlaw].
@@ 5. DRAWBACKS AND FUTURE DIRECTIONS OF LARGE LANGUAGE MODELS
Fairness: LLMs are widely employed for decision-making across various domains, impacting individuals and society. However, these decisions might exhibit bias against marginalized populations. For instance, when utilized in screening resumes for programming positions, LLMs might exhibit a bias favoring male candidates, demonstrating evident gender discrimination. This bias stems from the training of LLMs on extensive and unstructured Internet data, where they inherit stereotypes, misrepresentations, derogatory language, and exclusionary behaviors. These aspects disproportionately affect already disadvantaged and marginalized groups [bender2021dangers, dodge2021documenting, sheng2021societal]. The commonly employed strategy to mitigate bias in LLMs is to remove biased data from the training dataset. However, this strategy does not entirely eliminate bias and often diminishes the effectiveness of language modeling [meade2021empirical]. [saxena2023missed] discuss that the root causes of these problems lie not only in technology but also in socially acceptable definitions of fairness and meaningful interventions to ensure the long-term well-being of all groups. Researchers should consider the needs of disadvantaged groups from the outset and design these technologies proactively, rather than simply reacting to the biases present in their designed systems [gallegos2023bias]. The documented bias in LLMs, as highlighted by studies like [blodgett2017racial, mei2023bias, mozafari2020hate], underscores the imperative for future efforts in this direction. Fairness: LLMs are widely employed for decision-making across various domains, impacting individuals and society. However, these decisions might exhibit bias against marginalized populations. For instance, when utilized in screening resumes for programming positions, LLMs might exhibit a bias favoring male candidates, demonstrating evident gender discrimination. This bias stems from the training of LLMs on extensive and unstructured Internet data, where they inherit stereotypes, misrepresentations, derogatory language, and exclusionary behaviors. These aspects disproportionately affect already disadvantaged and marginalized groups [bender2021dangers, dodge2021documenting, sheng2021societal]. The commonly employed strategy to mitigate bias in LLMs is to remove biased data from the training dataset. However, this strategy does not entirely eliminate bias and often diminishes the effectiveness of language modeling [meade2021empirical]. [saxena2023missed] discuss that the root causes of these problems lie not only in technology but also in socially acceptable definitions of fairness and meaningful interventions to ensure the long-term well-being of all groups. Researchers should consider the needs of disadvantaged groups from the outset and design these technologies proactively, rather than simply reacting to the biases present in their designed systems [gallegos2023bias]. The documented bias in LLMs, as highlighted by studies like [blodgett2017racial, mei2023bias], underscores the imperative for future efforts in this direction.
@@
Safety: LLMs have diverse applications in industries such as finance and healthcare. However, these applications also present significant security challenges. LLMs may generate outputs that diverge from the provided context, user input, or factual knowledge, a phenomenon referred to as “hallucination” [bang2023multitask, huang2023survey]. This undermines their reliability in real-world scenarios, emphasizing the vital necessity for precision, particularly in critical domains like medicine [dash2023evaluation, umapathi2023med]. As a primary way to avoid these problems, well-aligned LLMs can be developed by including humans in the training loop and using reinforcement learning from human feedback (RLHF) [ouyang2022training, gao2020pile]. To improve model safety, it is also important to include safety-related cues in the RLHF process, as shown in GPT-4 [achiam2023gpt]. However, RLHF relies heavily on high-quality human feedback data from professional annotators, which is costly and time-consuming. As a result, enhancing the RLHF framework is essential to alleviate the burden on human annotators. Additionally, exploring more efficient annotation methods with assured data quality, such as Large Language Models (LLMs), can assist in the annotation process and enhance overall safety.
@@
Intellectual Property: LLMs possess the capability to create content resembling human-generated material, potentially encroaching upon users’ intellectual property rights. These models rely on training and optimizing using users’ textual data, which may encompass personal information, proprietary knowledge, patented technology, and more. Notably, AI-generated code exemplifies this issue. Code-generation models trained on open-source repositories can produce programs resembling existing ones, potentially disregarding relevant licenses [yu2023codeipprompt]. For instance, legal action has been taken against Microsoft, GitHub, and OpenAI, alleging copyright infringement due to Copilot reproducing licensed code without adherence to licensing terms [^11] https://www.reuters.com/technology/microsoft-attracting-users-its-code-writing-generative-ai-software-2023-01-25/  [^12] https://www.reuters.com/legal/litigation/openai-microsoft-want-court-toss-lawsuit-accusing-them-abusing-open-source-code-2023-01-27/. Consequently, there’s a pressing need for heightened attention and regulation concerning intellectual property rights concerning LLMs. This is essential to safeguard the legitimate rights and interests of original creators, users, and the public while fostering and respecting the innovation and advancement of LLMs technology [li2023protecting].